OpenStudy (anonymous):

What is the equation of the line that passes through (5, -2) and (-3, 4)?

OpenStudy (anonymous):

okay the point slope form of a line is m(x1-x2)=y1-y2

OpenStudy (anonymous):

find the slope between these two points it will be m

OpenStudy (anonymous):

it will be -3/4=m=slope

OpenStudy (anonymous):

now subtitute the value of x into x2 and value of y into y2 and value of slope into m

OpenStudy (anonymous):

for x and y you can take the coordinates of any particuklar point given to you

OpenStudy (anonymous):

you will get -3/4(x1-5)=(y1+2)

OpenStudy (anonymous):

good luck bro keep up the good work!!!